solve for x.
4x(x-8)-5x(x-7)=-4

Answer by ankor@dixie-net.com(22740)

solve for x.
4x(x-8)-5x(x-7)=-4
:
4x^2 - 32x - 5x^2 + 35x + 4 = 0
Combine like terms
4x^2 - 5x^2 - 32x + 35x + 4 = 0
-1x^2 + 3x + 4 = 0
Multiply by -1
x^2 - 3x - 4 = 0
Factors to
(x+1)(x-4) = 0
Two solutions
x = -1
x = 4
:
You can check the solutions in the original equation.